OpenStudy (anonymous):

find d/dx of f(x) = arcsin(x+1) how would i approach this

OpenStudy (anonymous):

just plug in the formula?

OpenStudy (anonymous):

yes, just use the formula for the derivative

OpenStudy (anonymous):

ty

OpenStudy (phi):

There is a very useful theorem that says that if you have y= f(x) and the inverse function x= g(y) then g'(y)= 1/f'(x) here y = asin(x-1) let u= x-1 y= asin(u) u= sin(y) du = cos(y) dy so dy/du = 1/cos(y) use sin^2(y)+cos^2(y)= 1 to rewrite cos(y) as sqrt(1- sin^2(y)) dy/du= 1/sqrt(1-sin^2(y)) but sin(y)= u, so this is dy/du = 1/sqrt(1-u^2) replace u with x-1 to finish this off